OpenStudy (zuha):

The table shows the height of a plant as it grows. Which equation in point-slope form gives the plant’s height at any time? Time (months) Plant Height (cm) 3 21 5 35 7 49 9 63 Choose the correct answer: y – 21 = 7/2(x – 3) y – 21 = 7(x – 3) y – 3 = 7/2(x – 21) The relationship is nonlinear.

OpenStudy (anonymous):

2nd choice bro since the slope if you take the first ex and the last

OpenStudy (anonymous):

you get 63-21/9-6 which is 42/6 =7

OpenStudy (zpupster):

he meant 63-21/9-3 which is 42/6 =7 so m =7 the slope in the equation y-y1=m(x-x1)

OpenStudy (zuha):

thanks both of you :)